As part of a flight training, a student pilot was required to fly to an airport and then return. the average speed on the way to the airport was 125 mph, and the average speed returning was 218.75 mph. Find the distance between the two airports if the total flying time was 8.25 hours.
1
Expert's answer
2012-07-17T09:38:51-0400
d = 125 * t1 = 218.75 * t2
t1 + t2 = 8.25, so t1 = 8.25 - t2. Getting:
125
* (8.25 - t2) = 218.75 * t2, 1031.25 - 125 * t2 = 218.75 * t2,
(218.75 + 125)
* t2 = 1031.25, t2 = 3 h
d = 218.75 * t2 = 218.75 * 3 = 656.25 miles